OpenStudy (anonymous):

The slope of the line that passes through the origin and the point (1, -3) is?

OpenStudy (anonymous):

y+3=-3/1(x-1) y+3=-3x+1 3x+y+2=0

OpenStudy (anonymous):

a. 1/3 b.-3 c.-1/3 d.3

OpenStudy (anonymous):

no slope is y-y1/x-x1 (-3-0)/(1-0) -3 is slope

OpenStudy (anonymous):

-3 is the answer?

OpenStudy (anonymous):

yaa acc 2 me